40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 149 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  9371   Wed Nov 13 01:35:40 2013 JenneUpdateLSCPRM motion causing trouble?

Quote:

* Still need to finish calculating what could be causing our big arm power fluctuations (Test mass angular motion?  PRM angular motion? ALS noise?) (Calculation)

 I think that our problem of seeing significant arm power fluctuations while we bring the arms into resonance during PRMI+arms tests is coming from PRM motion. I've done 3 calculations, so I will describe below why I think the first two are not the culprit, and then why I think the PRM motion is our dominant problem.

===============================================================

ALS length fluctuations

Arm length fluctuations seem not to be a huge problem for us right now, in terms of what is causing our arm power fluctuations.

What I have done is to calculate the derivative of the power in the arm cavity, using the power buildup that optickle gives me. The interferometer configuration I'm using is PRFPMI, and I'm doing a CARM sweep. Then, I look at the power in one arm cavity. The derivative gives me Watts buildup per meter CARM motion, at various CARM offsets. Then, I multiply the derivative by 60 nm, which is my memory of the latest good rms motion of the ALS system here at the 40m. I finally divide by the carrier buildup in the arm at each offset, to give me an approximation of the RIN at any CARM offset.

I don't know exactly what the calibration is for our ALS offset counts, but since we are not seeing maximum arm cavity buildup yet, we aren't very close to zero CARM offset. 

From this plot, I conclude that we have to be quite close to zero offset for arm length fluctuations to explain the large arm power fluctuations we have been seeing.

 xarmRIN_vs_CARM.png

=======================================================================

AS port contrast defect from ETM motion

For this calculation, I considered how much AS port contrast defect we might expect to see given some ETM motion. From that, I considered what the effect would be on the power recycling buildup. 

Rather than doing the integrals out, I ended up doing a numerical analysis. I created 2 Gaussian beams, subtracted the fields, then calculated the total power left. I did this for several separations of the beams to get a plot of contrast defect vs. separation. My simulated Gaussian beams have a FWHM of 1 unit, so the x-axis of the plot below is in units of spot motion normalized by spot size. 

Unfortunately, my normalization isn't perfect, so 2 perfectly constructively interfering beams have a total power of 0.3, so my y-axis should all be divided by 0.3. 

The actual beam separation that we might expect at the AS port from some ETM motion (of order 1e-6 radians) causing some beam axis shift is of the order 1e-5 meters, while the beam spot size is of the order 1e-3 meters. So, in normalized units, that's about 1e-2. I probably should change the x-axis to log as well, but you can see that the contrast defect for that size beam separation is very small. To make a significant difference in the power recycling cavity gain, the contrast defect, which is the Michelson transmission, should be close to the transmission of the PRM. Since that's not true, I conclude that ETM angular motion leading to PRC losses is not an issue. 

I still haven't calculated the effect of ITM motion, nor have I calculated either test mass' angular effect directly on arm cavity power loss, so those are yet to be done, although I suspect that they aren't our problem either. 

ASportOverlap_moveGaussianBeams.png

========================================================================

PRM motion

 

I think that the PRM moving around, thus causing a loss in recycling gain, is our major problem. 

First, how do I conclude that, then some thoughts on why the PRM is moving at all. 

=========

theta = 12e-6 radians (ref: oplev plot from elog 9338 last week)

L = 6.781 meters

g = 0.94

a = theta * L /(1-g) = 0.0014 meters axis displacement

w0 = 3e-3 meters = spot size at ITM

a^2/w0^2 = 0.204 ==>> 20% power loss into higher order modes due to PRM motion.

That means 20% less power circulating, hitting the ITMs, so less power going into the arm cavities, so less power buildup. This isn't 50%, but it is fairly substantial, using angular fluctuation numbers that we saw during our PRMI+arms test last week. If you look at the oplev plot from that test, you will notice that when the arm power is high (as is POP), the PRM moves significantly more than when the carrier buildup in the cavities was low. The rms motion is not 12 urad, but the peak-to-peak motion can occasionally be that large. 

So, why is that? Rana and I had a look, and it is clear that there is a difference in PRM motion when the IFO is aligned and flashing, versus aligned, but PSL shutter is closed. Written the cavities flash, the PRM gets a kick. Our current theory is that some scattered light in the PRC or the BS chamber is getting into the PRM's OSEMs, causing a spike in their error signal, and this causes the damping loops to push on the optic. 

We should think a little more on why the PRM is moving so much more that any other optic while the power is building up, and if there is anything we can do about the situation without venting. If we have to, we should consider putting aluminum foil beam blocks to protect the PRM's OSEMs. 

  9370   Tue Nov 12 23:48:23 2013 RANAUpdateIOOPSL pointing monitoring

Since I saw that the trend was good, I aligned the MC refl path to the existing IMC alignment:

  1. removed a broken IRIS that was clipping the reflected beam (and its mount)
  2. moved the first 1" diameter steering mirror on the high power path after the 2" diameter R=10% steering mirror. It was not centered.
  3. Moved the lens just upstream of the LSC RFPD away from the PD by ~5 mm. The beam going towards the WFS was too close to this mount and I could see some glow.
  4. Centered the beam on all optics in the WFS path and then the WFS DC.
  5. Centered beam on LSC RFPD.

The reflected spots from the PD are not hitting the dump correctly. WE need to machine a shorter post to lower the dump by ~1 cm to catch the beams.

  9369   Tue Nov 12 23:05:06 2013 ranaUpdatePEMthump noise from particle counter

Quote:
The particle counter is back on the IOOC location on a piece of FOAM
It needs this isolation, so when the pump is running, it's not shaking things.
The counter was counting for 6 sec and it was on holding for 20 mins.
Now I set the counter for 20 sec so it is easy to recognise it's signal and it holds for 2min only.
This will set the alarm handler in action.

Atm1: 40 mins plot
PEM-ACC_MC2_x,y,z up to 13 mins: pcounter at MC2 table, clamped, counting for 20s and holds for 2 mins
PEM-ACC_MC2_x,y,z from 13 to 26 mins: pcounter at MC2 table, not clamped, seated on 2" foam, counting 20s and holds for 2 mins
PEM-ACC_MC1_x,y,z from 26 to 40 mins: pcounter at MC1_IOOC location, not clamped, seated on 2" foam, counting 20s and holds for 2 mins

Rana won the bet

This noise source from 2008 is back. Somehow the particle counter snuck back up to the top of the BS chamber. Bad, noisy particle counter. Let's move it away from out sensitive optics and put it on the wall next to the router or maybe on some toolbox. Not on optics tables, chambers, or near them.
  9368   Tue Nov 12 21:50:01 2013 ranaUpdate zita network configured

 

 I installed 'nfs-client' on zita (the StripTool terminal). It now has mounted all the shared disks, but still can't do StripTool since its a 32-bit machine and our StripTool is 64.

  9367   Tue Nov 12 16:49:22 2013 JenneUpdateLSCXend QPD and Whitening board pulled

Quote:

* Whitening for the transmission QPDs needs to be thought about more carefully. (Calculation, then hardware)

 I have the X end transmission QPD, as well as the whitening board, out on the electronics bench.  Since the Thorlabs high-gain TRX PD also goes through this whitening board, we have no transmission signal for the Xarm at this time. The whitening board was in the left-most slot, of the top crate in the Xend rack.  The only cables that exist for it (like the Yend), are the ribbon from the QPD, the 4-pin lemo from the Thorlabs PD, and the ribbon going to the ADC.

I have taken photos, and want to make sure that I know what is going on on the circuits, before I put them back in. 

The QPD:

IMG_1262.JPG

IMG_1260.JPG

The whitening board:

IMG_1264.JPG

IMG_1276.JPG

  9366   Tue Nov 12 15:04:35 2013 ranaUpdateCDSFE Web view was fixed

Quote:

 Seems partially broken again. Not updating for most of the FE. I've commented out the cron lines for this as well as the mostly broken MEDM Snapshots job. I'm in the process of adding them to the megatron cron (since that machine is at least running 64 bit Ubuntu 12, instead of 32-bit CentOS)

 https://nodus.ligo.caltech.edu:30889/medm/screenshot.html

Seems to now be working. I made several fixes to the scripts to get it working again:

  1. changed TCSH scripts to BASH. Used /usr/bin/env to find bash.
  2. fixed stdout and stderr redirection so that we could see all error messages.
  3. made the PERL scripts executable. most of the PERL errors are not being logged yet.
  4. fixed paths for the MEDM screens to point to the right directories.
  5. the screen cap only works on screens which pop open on the left monitor, so I edited the screens so that they open up there by default.
  6. moved the CRON jobs from mafalda over to megatron. Mafalda no longer is running any crons.
  7. op540m used to run the 3 projector StripTool displays and have its screen dumped for this web page. Now zita is doing it, but I don't know how to make zita dump her screen.
  9365   Mon Nov 11 22:35:45 2013 RANAUpdateIOOPSL pointing monitoring

Since the pointing has gone bad again, I went to the PSL to investigate. Found some bad things and removed them:

1) There was a stopped down iris AGAIN in the main beam path, after the newly installed mirror mount. I opened it. Stop closing irises in the beam path.

2) The beam dump for the IOO QPD reflection was just some black aluminum. That is not a real dump. I removed it. We need two razor blade dumps for this.

3) There was an ND filter wheel (???) after one of the PMC steering mirrors. This is not good noise / optics practice. I removed it and dumped the beam in a real dump. No elog about this ?!#?

 

The attached trend shows the last 20 days. The big step ~2 weeks ago is when Steve replaced the steering mirror mount with the steel one. I don't understand the drift that comes after that.

 

Today I also spent ~1 hour repairing the Aldabella laptop. Whoever moved it from the PSL area to the SP table seems to have corrupted the disk by improper shutdown. Please stop shutting the lid and disconnecting it from the AC power unless you want to be fixing it. Its now running in some recovery mode. Lets leave it where it is next to the PSL and MC1.

I steered the MC suspensions back to where they were on the trends before the PSL mirror mount swap and then aligned the PSL beam into it by touching the last 2 steel mounts. Once the alignment was good without WFS, I centered the beams on the IOO QPDs. If it behaves good overnight, I will center the unlocked beams on the MC WFS.

 

Please stay off the PSL for a couple days if you can so that we can watch the drift. This means no opening the doors, turning on the lights, or heavy work around there.

Attachment 1: qpd.pdf
qpd.pdf
  9364   Mon Nov 11 12:19:36 2013 ranaUpdateCDSFE Web view was fixed

Quote:

FE Web view was broken for a long time. It was fixed now.

The problem was that path names were not fixed when we moved the models from the old local place to the SVN structure.

The auto updating script (/cvs/cds/rtcds/caltech/c1/scripts/AutoUpdate/update_webview.cron) is running on Mafalda.

Link to the web view: https://nodus.ligo.caltech.edu:30889/FE/

 Seems partially broken again. Not updating for most of the FE. I've commented out the cron lines for this as well as the mostly broken MEDM Snapshots job. I'm in the process of adding them to the megatron cron (since that machine is at least running 64 bit Ubuntu 12, instead of 32-bit CentOS)

  9363   Sat Nov 9 10:25:43 2013 KojiUpdateLSCNew Broadband PD for POP 22/110

General Remarks on the BBPD

- To form the LC network: Use fixed SMD inductors from Coilcraft. SMD tunable capacitors are found in the shelf right next to Steve's desk.
  If the tuning is too coarse, combine an appropriate fixed ceramic SMC C and the tunable C (in parallel, of course)

- L1/C1a/C1b pads are specifically designed for an additional notch

- Another notch at the diode stage can be formed between the middle PD pin (just left of the marking "C3b") to the large GND pad (between C1a/C1b to C3a).
  You have to scratch off the green resin with a small flat screw driver (or anything similar)

- A notch at the amplifier stage can be formed between the output of MAR-6SM ("+" marking)  and one of the GND pads (left side of the "U1" marking)

- The original design of the PD is broadband. So additional notches on the diode stage provides notches and resonances.
  Check if the resonances do not hit the signal frequencies.

- One would think the PD can have resonant feature to reduce the coupling of the undesired signals.
  In some sense it is possible but it will be different from the usual resonant tank circuit in the following two points.

  * Just adding a parallel L between the cathode and ground does not work. As this DC current should be directed to the DC path,
    L&C combo should be added. In fact this actually give a notch-resonance pair. This C should be big enough so that you can ignore it
    at the target resonant frequency. Supply complimentary small C if necessary to keep low impedance of the Cs at the target frequency.
    (i.e. Check SRF - self-resonant frequency of the big C)

  * Since the input impedance of MAR-6SM is 50Ohm, the top of the resonant curve will be cut at 50Ohm. So the resultant shape looks
    like a bandpass rather than a resonance.

- So in total, simulation of the circuit is very important to shape the transimpedance. And, consider the circuit can not be formed as simulated
  because of many practical imperfections like stray Ls and Cs.

 

  9362   Fri Nov 8 18:12:21 2013 JenneUpdateLSCPRFPMI: Not crossing any resonances

Quote:

There are several things at this point that we know we need to look into:

* Simulate an arm sweep, up to many orders of the sidebands, to see how close to the carrier resonance any sideband resonances might be.  If something like the 4th order sideband resonates, and then beats with a 1st order sideband, is that signal big enough to disturb our 3f locking of the PRMI / DRMI?  We want to be holding the arms off resonance with ALS closer to the carrier than any "important" sideband resonances (where the definition of "important" is still undetermined).  (Simulation)

 I have done a sweep of CARM, while looking at the fields inside of one arm (I've chosen the Xarm), to see where any resonances might be, that could be causing us trouble in keeping the PRMI locked as we bring the arms into resonance. 

2f_resonances_Xarm_CARMsweep.png

Since Gabriele pointed out to me that we're using the 3x55MHz signal for locking, we should be most concerned about resonances of the higher orders of 55, and not of 11.  So, on this plot, I have up to the 6th order 55 MHz sidebands, which are 332 MHz.  Although the Matlab default color chart has wrapped around, it's clear that the carrier is the carrier, and the +4f2, which is the same blue, is not the giant central peak.  So, it's kind of clear which trace is which, even though the legend colors are degenerate.  Also, the main point that I want to show here is that there is nothing going on near the carrier, with any relevant amplitude.  The nearest things are the plus and minus 55 MHz sidebands themselves, and they're more than 50 nm away from the carrier. 

Recalling from elog 9122, the PRFPMI and DRFPMI linewidths are about 40pm.  50pm away from the resonant point is ~1/10 the power, and 100pm away from the resonant point is ~1/100 the power.  So, 50 nm is a looooong ways away. 

Just for kicks, here is a plot of all the resonances of the 1f and 2f modulation frequencies, up to 30*f1, which is the same 6*f2:

AllModFreq_resonances_Xarm_CARMsweep.png

The resonances which are "close" to the carrier are the 9th order 11 MHz sidebands, and they're 280pm from the carrier, so twice as far as we need to be, to get our arm powers to ~1/100 of the maximum, and, they're a factor of ~1e4 smaller than the carrier.

  9361   Fri Nov 8 17:19:27 2013 JenneUpdateLSCNew Broadband PD for POP 22/110

Here is a photo of the board inside the broadband photodiode (one of them) that I took from the Gyro experiment:

IMG_1255.JPG

This PD is Serial Number S1200271.

We need to have a look at the schematic, figure out what's in here now, and then modify this to be useful (appropriate resonances / notches, as well as amplification) for POP 22/110.

  9360   Thu Nov 7 14:39:49 2013 JenneUpdateLSCLock Acquisition Game Plan

Between the 40m meeting, and chatting with Gabriele, there was lots of talking yesterday about our 40m Lock Acquisition game plan. 

From those talks, here is my current understanding of the plan, in a Ward-style cartoon:

 LockAcquisition_7Nov2013.pdf(This is a 2 page document - description of steps is on 2nd page)

If you look closely, you will notice that there are several places that I have used "?" rather than numbers, to indicate what RFPD signal we should be using.  To fill these in, I need to look at some more simulations, and think more carefully about what signals exist at what ports, and what SNR we have at each of those ports.

Also, while the overall scale of the arm power plot is correct, the power level at each step is totally arbitrary right now, and should just be taken to mean places (in time) where the CARM offset is reduced a little more.


There are several things at this point that we know we need to look into:

* POP 22/110 PD and filtering electronics should be switched to a broadband PD, rather than the Thorlabs PD + Miniciruits filters. (Hardware)

* Whitening for the transmission QPDs needs to be thought about more carefully. (Calculation, then hardware)

* Chose a good SNR REFL DC signal, which may or may not be from the PD we are currently using (I think it's the DC of REFL11, but I'll have to check). (Calculation)

* For DRMI locking, what is the size of the SRCL error signal at AS55, AS165, and the REFL ports?  Do we need to lock with AS port, and then switch over to a REFL 3f port, to make acquisition easier?  (Simulation)

* Similarly, I want to make the equivalent of Figure 3 of T1000294, with our 40m parameters. (Simulation)

* To set the phase of AS110, simulate the demod phase of AS110 in both DRMI and SRMI cases.  If no (significant) change, maybe we can set the phase in the real system by misaligning the PRM, and watching the SRMI flash.  (Simulation)

* Simulate an arm sweep, up to many orders of the sidebands, to see how close to the carrier resonance any sideband resonances might be.  If something like the 4th order sideband resonates, and then beats with a 1st order sideband, is that signal big enough to disturb our 3f locking of the PRMI / DRMI?  We want to be holding the arms off resonance with ALS closer to the carrier than any "important" sideband resonances (where the definition of "important" is still undetermined).  (Simulation)

* Check if we can hand DARM from the DC transmission signals to the final RF signal while we still have a large CARM offset. Is there a point where the CARM offset is too large, and we must be still using the DC signals? (Simulation)

* At what arm power level can we transition from ALS to IR DC transmission signals for the individual arms? (Simulation)

* Still need to finish calculating what could be causing our big arm power fluctuations (Test mass angular motion?  PRM angular motion? ALS noise?) (Calculation)


Replys, and comments are welcome, particularly to help me understand where I may have (likely did) go wrong in drawing my cartoon. 

  9359   Thu Nov 7 11:19:04 2013 SteveUpdateVACcrane work POSTPONED to Monday

Quote:

 

The smoke alarms were turned off and surrounding areas were covered with plastic.

The folding I-beam was ground down to be in level with the main beam.

Load bearing cable moved into correct position. New folding spring installed.

Crane calibration was done at 500 lbs at the end of the fully extended jib.

Than we realized that the rotating wheel limit switch stopped working.

This means that the crane is still out of order.

 

 New limit switch will be installed tomorrow morning

Konecranes postponed installation Friday morning Nov. 8

Friday 5pm : Konacranes promising to be here 8am Monday, Nov 11

It was rescheduled on Tuesday again for Wednesday, Nov 13

  9358   Thu Nov 7 08:57:20 2013 SteveUpdateIOOPSL pointing monitoring

 The qpd sees the power drop as position change.

The laser  monitoring screen shows little changes of the Innolight 2W output. See elog 9292 to compare

So why does the PMC downgrade if the laser output is stationary ?

The PMC-T power is down to 0.75V  The auto locker does maximize power output.

It needs a manual alignment touch up.

 

 

Attachment 1: PMC11072013.jpg
PMC11072013.jpg
Attachment 2: PMCTqpdMon.png
PMCTqpdMon.png
  9357   Wed Nov 6 17:21:58 2013 JamitUpdateCDSFB not talking to LSC?

Quote:

Something funny is going on with the framebuilder's communication with the LSC machine. 

This is a different failure mode / error than I have seen before.  It's not the type of problem that is solved by restarting the mxstreams (that is indicated by also the 2 blocks on top of one another, that are green on the lsc machine right now, being red), although I did try that, before I looked closer and realized that that wasn't the problem.

ssh-ing to c1lsc, and doing a "rtcds restart all" seems to be fixing the problem.  Both c1oaf and c1cal needed another round of restarting, because they needed their BURT buttons pressed manually.  All of the models on the lsc machine are running fine now, though.

Here's a screenshot of the CDS overview screen, with the error lights:

Screenshot-Untitled_Window-1.png

This definitely looks like a timing problem on the c1lsc front end computer.  The red lights on the left mean that the timing synchronization is lost at the user model.  I'm perplexed why it looks like the IOP is not seeing the same error, though, since it should originate at the ADC.  The red lights to the right just mean the timing synchronization is lost with the DAQ, which is too be expected given a timing loss at the front end.

We'll have to take a closer look when this happens again.

  9356   Wed Nov 6 15:59:41 2013 manasaUpdateElectronicsIOO rack +/-5V power supplies

Quote:

The power supply to the ADC box on the IOO rack (that reads the beat I & Q signals) was pulled out because it did not run through any fuse and was connected directly to the power supply. 

There were already connections running from the +/-5 V power supply. They were powering the mode cleaner demod board rack. In order to remove the ADC power connectors from the power supply, I notified Jenne in the control room because turning off the power supply would affect the MC. I switched off the +/-5V power supplies at the same time. The ADC power connectors were removed. The +/-5V power supplies were then turned ON again at the same time. Jenne relocked the MC after this.

I have still not connected the ADC to the fuse rack power supply because this requires the +/-5V power supplies to be turned OFF again in order to pull out new connections from the fuse rack and I need to make a new ADC power connector with thicker wires.

I switched OFF the +/-5V power supplies on the IOO rack to hook up the ADC power connectors through 250mA fuses to +/-5V. Since these power supplies were powering the MC demod boards, MC remained unlocked during the process. I turned the power supplies back ON and MC relocked itself after this.

  9355   Wed Nov 6 15:57:22 2013 JenneUpdateLSCPower Supply solution

We have decided that, rather than replacing the power source for the amplifiers that are on the rack, and leaving the Thorlabs PD as POP22/110, we will remove all of the temporary elements, and put in something more permanent.

So, I have taken the broadband PDs from Zach's Gyro experiment in the ATF.  We will figure out what needs to be done to modify these to notch out unwanted frequencies, and amplify the signal nicely.  We will also create a pair of cables - one for power from the LSC rack, and one for signal back to the LSC rack.  Then we'll swap out the currently installed Thorlabs PD and replace it with a broadband PD.

  9354   Wed Nov 6 15:12:01 2013 JenneUpdateCDSFB not talking to LSC?

Something funny is going on with the framebuilder's communication with the LSC machine. 

This is a different failure mode / error than I have seen before.  It's not the type of problem that is solved by restarting the mxstreams (that is indicated by also the 2 blocks on top of one another, that are green on the lsc machine right now, being red), although I did try that, before I looked closer and realized that that wasn't the problem.

ssh-ing to c1lsc, and doing a "rtcds restart all" seems to be fixing the problem.  Both c1oaf and c1cal needed another round of restarting, because they needed their BURT buttons pressed manually.  All of the models on the lsc machine are running fine now, though.

Here's a screenshot of the CDS overview screen, with the error lights:

Screenshot-Untitled_Window-1.png

  9353   Wed Nov 6 14:47:41 2013 SteveFrogsLSCDin connectors added at 1Y2

 The north side of the LSC rack is full. I installed more DIN connectors with fuses on the south side of the rack 1Y2

The access to this may be a little bit awkward. You just remove the connector, wire it and put it back in.

 

Attachment 1: 1Y2moreDinFuses.jpg
1Y2moreDinFuses.jpg
  9352   Wed Nov 6 08:33:53 2013 SteveUpdateIOOPoiting changes of PSL

Quote:

 Full list tomorrow: IP-Ang & Pos, ETMY-T, ETMY-Oplev, ETMX-T, IOO-Ang & Pos

 RA: No one in the control room this evening can understand what this ELOG means. Please use more words.

 Yesterday the last steering mirror mount on the PSL was changed, Manasa recovered the MC alignment and Jenne locked the arms.

 I centered the following qpds:  ASC-IBQPD, LSC-TRY, SUS-ETMY_OPLEV, LSC-TRX, SUS_ETMX_OPLEV

 Touching the PSL pointing IOO-QPD_ANG & POS was a mistake. We lost the reference of the well refined MC input.

 

One and 20 days TRENDS  plot showing the PSL output drift in pitch can be power drop

However initial pointing is amazingly good. ( I wonder about the lens in front of the qpd ?)

Attachment 1: 1dayTRENDofPOITING.png
1dayTRENDofPOITING.png
Attachment 2: 20dayTRENDioo.png
20dayTRENDioo.png
  9351   Tue Nov 5 19:55:12 2013 ranaUpdateSUSoplev XY-plots reflect new calibration

I used the same OSEM SUSPIT/YAW method as before to calibrate the ETMY optical lever signals. They were off by a factor of ~10.

ETMY Pitch     300  /  26    (old/new)     urad/counts

ETMY Yaw       300  /  31    (old/new)     urad/counts

These should be redone with the Kakeru / Ottaway arm cavity power technique if we want to get better than ~30% accuracy.

  9350   Tue Nov 5 19:50:07 2013 manasaUpdateElectronicsIOO rack +/-5V power supplies

The power supply to the ADC box on the IOO rack (that reads the beat I & Q signals) was pulled out because it did not run through any fuse and was connected directly to the power supply. 

There were already connections running from the +/-5 V power supply. They were powering the mode cleaner demod board rack. In order to remove the ADC power connectors from the power supply, I notified Jenne in the control room because turning off the power supply would affect the MC. I switched off the +/-5V power supplies at the same time. The ADC power connectors were removed. The +/-5V power supplies were then turned ON again at the same time. Jenne relocked the MC after this.

I have still not connected the ADC to the fuse rack power supply because this requires the +/-5V power supplies to be turned OFF again in order to pull out new connections from the fuse rack and I need to make a new ADC power connector with thicker wires.

  9349   Tue Nov 5 19:39:27 2013 JenneUpdateLSCOpLev time series

[Rana, Jenne]

We looked at the time series for all the oplevs except the BS, from last Tuesday night, during a time when we were building up the power in the arms.  We conclude from a 400 second stretch of data that there is not discernible difference in the amount of motion of any optic, when the cavities are at medium power, and when they're at low power.  Note however, that we don't have such a nice stretch of data for the really high powers, so the maximum arm power in these plots is around 5.  Both the TRX and TRY signals look fairly stationary up to powers of 1 or 2, but once you get to 4 or 5, the power fluctuations are much more significant.  So, since this isn't caused by any optic moving more, perhaps it's just that we're more sensitive to optic motion when we're closer to resonance in the arms.

However, from this plot, it looks like the ETMY is moving much more than any other optic.  On the other hand, ETMY has not ever been calibrated (there's an arbitrary 300 in there for the calibration numbers on the ETMY oplev screen).  So, perhaps it's not actually moving any more than other optics.  We should calibrate the ETM oplevs nicely, so we have some real numbers in there.  ETMX also only is roughly calibrated, relative to the OSEMs.  We should either do the move-the-QPD calibration, or a Kakeru-style pitch and yaw some mirrors and look at transmitted power.

Traces on this xml file have been filtered with DTT, using zpk([0],[0.03],1,"n").

OpLevs_during_PRMI_2arms.pdf

 

  9348   Tue Nov 5 17:12:48 2013 JenneFrogsLSCillegal power supply about to expire

Quote:

 Is this your illegally installed HP bench power supply?

 Steve has promised to add another row of fuses to the LSC rack first thing in the morning.  Then, during Wednesday Chores, we can move the wires from the power supply to the fused power.

STEVE:  NEVER MIND about doing this in the morning.  Let's chat at the lunch meeting about what needs to be done to power things down, then back up again, in a nice order, and we can do it after lunch.  

So, please do not do anything to the LSC rack tomorrow!  Thank you.

  9347   Tue Nov 5 17:12:34 2013 SteveUpdateLSCcentered qpds

 Full list tomorrow: IP-Ang & Pos, ETMY-T, ETMY-Oplev, ETMX-T, IOO-Ang & Pos

 RA: No one in the control room this evening can understand what this ELOG means. Please use more words.

Attachment 1: recentered.png
recentered.png
  9346   Tue Nov 5 16:47:19 2013 ranaFrogsLSCillegal power supply about to expire

 Is this your illegally installed HP bench power supply?

20131023_222351.jpg

If so, or if not but you care about the signal that passes through these amplifiers, I suggest you remove this temporary power supply and wire the power from the rack power supplies through the fuse blocks and possibly use a voltage regulator.

In 24 hours, that power supply will be disconnected and the wires snipped if they are still there.

  9345   Tue Nov 5 16:47:09 2013 JenneUpdateLSCEnd transmission QPDs

I think Steve is trying to align the end transmission QPDs, since the arms are locked nicely right now.  I noticed that the QPDX pitch and yaw signals were digital zeros.  A quick look determined that the QPDX matrix to go from 4 quadrants to 3 degrees of freedom had been filled in for the POS row, but not pitch and yaw.  So, I copied the QPDY matrix over to QPDX (so the ordering of the rows and columns is assumed to be the same). 

Hopefully this will get us close to centered, but I suppose we ought to check really which quadrant is which, by shining a laser pointer at each quad at each end.

  9344   Tue Nov 5 16:39:54 2013 GabrieleUpdateLSCThoughts on Transition to IR

Quote:

Gabriele and I talked for a while on Wednesday afternoon about ideas for transitioning to IR control, from ALS. 

I think one of the baseline ideas was to use the sqrt(transmission) as an error signal.  Gabriele pointed out to me that to have a linear signal, really what we need is sqrt( [max transmission] - [current transmission] ), and this requires good knowledge of the maximum transmission that we expect.  However, we can't really measure this max transmission, since we aren't yet able to hold the arms that close to resonance.  If we get this number wrong, the error signal close to the resonance won't be very good.

Gabriele suggested maybe using just the raw transmission signal.  When we're near the half-resonance point, the transmission gives us an approximately linear signal, although it becomes totally non-linear as we get close to resonance.  Using this technique, however, requires lowering the finesse of PRCL by putting in a medium-large MICH offset, so that the PRC is lossy.  This lowering of the PRC finesse prevents the coupled-cavity linewidth of the arm to get too tiny.  Apparently this trick was very handy for Virgo when locking the PRFPMI, but it's not so clear that it will work for the DRFPMI, because the signal recycling cavity complicates things.

I need to look at, and meditate over, some Optickle simulations before I say much else about this stuff.

 The idea of introducing a large MICH offset to reduce the PRC finesse might help us to get rid of the transmitted power signal. We might be able to increase enough the line width of the double cavity to make it larger than the ASL length fluctuations. Then we can switch from ASL to the IR demodulated signal without transitioning through the power signal.

  9343   Tue Nov 5 08:44:21 2013 SteveUpdateIOOafter last steering mirror mount swap

 The IOO Angle and IOO Position qpds  were recentered after this entry.

 Suggested corrections in elog entry #9323 are completed:

 1,  last steering mirror mount replaced by Polanski mount

 2,  PSL output shutter mount reconfigured

 IOO qpds are not centered. I failed to connect laptops to 40MARSian network.

Attachment 1: after_Shutter_SMm_installed.png
after_Shutter_SMm_installed.png
Attachment 2: afterlastSMmountswap.jpg
afterlastSMmountswap.jpg
  9342   Tue Nov 5 00:39:43 2013 manasaUpdateIOOIFO alignment tuning

Information acknowledged from Steve:

The last steering mirror mount for IR on the PSL was swapped for a more robust one. Prior to swapping the ibeam positions on the PSL IOO QPDS in ang and pos were recorded.

What I did henceforth:

1. Once the last steering mirror was installed, I walked the beam to restore input pointing using the last 2 steering mirrors. It needed a lot of work in yaw as expected.

2. When the input pointing was recovered, MC locked right away in TEM00. I measured the MC spot positions and compared it with Jenne's measurement made prior to the swap. The spot positions were pretty close.

3. The input pointing was adjusted in pitch and yaw (on the last steering mirror) in small steps. MC alignment was recovered and spot positions were measured each time. After several iterations, the MC spot positions were pretty much centered. I recentered the WFS and reset the WFS offsets. MC is now locked with WFS enabled at ~16900 counts.

MC_spot.png

4. Since the arms were aligned this morning, I used the Y arm as reference and corrected for the input pointing using tip-tilts.

5. Arms locked right away. Note: ASS doesn't seem to be doing it's job. I had to manually align the arms for maximum on TRX and TRY.

6. MICH and PRMI lock were also recovered.

7. I started to check the status with ALS as well. But for reasons unknown, I don't see any ADC counts corresponding to the beat note. Looking at the beatbox there aren't any signs of disconnected cables.  I am saving this as a morning job to fix it.

  9341   Mon Nov 4 23:11:00 2013 JenneUpdateLSCSmall updates to LSC screen

I made some small edits to the LSC screen. 

* When I added columns for the new AS110 PD, I had forgotten to make the Trigger matrix and Power Normalization matrix icons on the screen bigger, so we weren't seeing the last 2 columns in the overview screen.

* I added "show if not zero" oscillator icons to the Sensing Matrix part of the LSC overview screen, so that it's easier at a glance to see that there is an oscillator on.

  9340   Mon Nov 4 18:24:15 2013 KojiUpdateLSCThoughts on Transition to IR

 You have the data. Why don't you just calculate 1/SQRT(TRX)?

...yeah, you can calculate it but of course you don't have no any reference for the true displacement...

  9339   Mon Nov 4 17:08:23 2013 JenneUpdateLSCThoughts on Transition to IR

Gabriele and I talked for a while on Wednesday afternoon about ideas for transitioning to IR control, from ALS. 

I think one of the baseline ideas was to use the sqrt(transmission) as an error signal.  Gabriele pointed out to me that to have a linear signal, really what we need is sqrt( [max transmission] - [current transmission] ), and this requires good knowledge of the maximum transmission that we expect.  However, we can't really measure this max transmission, since we aren't yet able to hold the arms that close to resonance.  If we get this number wrong, the error signal close to the resonance won't be very good.

Gabriele suggested maybe using just the raw transmission signal.  When we're near the half-resonance point, the transmission gives us an approximately linear signal, although it becomes totally non-linear as we get close to resonance.  Using this technique, however, requires lowering the finesse of PRCL by putting in a medium-large MICH offset, so that the PRC is lossy.  This lowering of the PRC finesse prevents the coupled-cavity linewidth of the arm to get too tiny.  Apparently this trick was very handy for Virgo when locking the PRFPMI, but it's not so clear that it will work for the DRFPMI, because the signal recycling cavity complicates things.

I need to look at, and meditate over, some Optickle simulations before I say much else about this stuff.

  9338   Mon Nov 4 15:46:17 2013 JenneUpdateLSCThoughts and Conclusions from last week's PRMI+2arms attempt

5:31pm - This is still a work in progress, but I'm going to submit so that I save my writing so far. I think I'm done writing now.


First, a transcription of some of the notes that I took last Tuesday night, then a few looks at the data, and finally some thoughts on things to investigate.


MICH and PRCL Transfer Functions while arms brought in to resonance (both arms locked to ALS beatnotes):

This is summarized in elog 9317, which I made as we were finishing up Tuesday night.  Here's the full story though.  Note that I didn't save the data for these, I just took notes (and screenshots for the 1st TF).

POP22I was ~140 counts, POP110I was ~100 counts.

MICH gain = -2.0, PRCL gain = 0.070. 

First TF (used as reference for 2-10), PRMI locked on REFL165, Xarm transmission = 0.03, Yarm transmission = 0.05 (both arms off resonance).  MICH UGF~40Hz, PRCL UGF~80Hz.

MICH_40Hz.pngPRCL_80Hz.png

2: X=off-res (xarm not moved), Y=0.13, no change in TF

3: X=off-res (xarm not moved), Y=0.35, no change in TF

4: X=off-res (xarm not moved), Y=0.60, MICH high freq gain went up a little, otherwise no change (no change in either UGF)

5:  X=off-res (xarm not moved), Y=0.95, same as TF#4.

6: X=0.20, Y=1.10 (yarm not moved), same as TF#4

7:  X=0.40, Y=1.30 (yarm not moved), same as TF#4

8:  X=0.70, Y=1.55 (yarm not moved), same as TF#4

9: X=1.40, Y=2.20 (yarm not moved), same as TF#4

10: X=4.0, Y=4.0 (yarm not moved), PRCL UGF is 10Hz higher than TF#4, MICH UGF is 20Hz lower than TF#4.

11: (No TF taken), Xarm and Yarm transmission both around 20!  To get this, MICH FMs that were triggered, are no longer triggered to turn on.  Also, MICH gain was lowered to -0.15 and PRCL gain was increased to 0.1

12: (No TF taken), Xarm and Yarm transmissions both around 40!  The peaks could be higher, but we don't have the QPD ready yet.

After that, we started moving away from resonance, but we didn't take any more transfer functions.


OpLev spectra for different arm resonance values:

We were concerned that the ETMs and ITMs might be moving more, when the arms are resonating high power, due to some optical spring / radiation pressure effects, so I took spectra of oplevs at various arm transmissions.

I titled the first file "no lock", and unfortunately I don't remember what wasn't locked.  I think, however, that nothing at all was locked.  No PRMI, no arm ALS, no nothing.  Anyhow, here's the spectrum:

ALS_noLock.pdf

I have a measurement when the Yarm's transmission was 1, and the Xarm's transmission was 1.75.  This was a PRMI lock, with ALS holding the arms partially on resonance:

ALS_X1pt75Y1.pdf

Next up, I have a measurement when Yarm was 0.8, Xarm was 2.  Again, PRMI with the arms held by ALS:

ALS_X2Y0pt8.pdf

And finally, a measurement when Xarm was 5, Yarm was 4:

ALS_X5Y4.pdf

Just so we have a "real" reference, I have just now taken a set of oplev spectra, with the ITMs, ETMs and PRM restored, but I shut the PSL shutter, so there was no light flashing around pushing on things.  I noticed, when taking this data, that if the PSL shutter was open, so the PRFPMI is flashing (but LSC is off), the PRM oplev looks much like the original "no Lock" spectra, but when I closed the shutter, the oplev looks like the others.  So, perhaps when we're getting to really high powers, the PRM is getting pushed around a bit?

ALS_noLock_noLaser.pdf

Conclusions from OpLev Spectra:  At least up to these resonances (which is, admittedly, not that much), I do not see any difference in the oplev spectra at the different buildup power levels.  What I need to do is make sure to take oplev spectra next time we do the PRMI+2arms test when the arms are resonating a lot. 


Time series while bringing arms into resonance:

PRMI_2arms_29Oct2013_POPrin.png

I had wondered if, since the POP 22 and 110 values looked so shakey, we were increasing the PRCL RIN while we brought the arms into resonance.  You can see in the above time series that that's not true.  The left side of the plot is PRMI locked, arms held out of resonance using ALS.  First the Yarm is brought close to resonance, then the Xarm follows.  The RIN of the arms is maybe increasing a little bit as we get closer to resonance, but not by that much.  But there seems to be no correlation between arm power and RIN of the power recycling cavity.

Alternatively, here is some time series when the arm powers got pretty high:

PRMI_2arms_29Oct2013_POPrin_highArmPowers.png


Possible Saturation of Signals:

One possibility for our locklosses of PRMI is that some signal somewhere is saturating, so here are some plots showing that that's not true for the error and control signals for the PRMI:

PRMI_2arms_29Oct2013_LSCcontrolSignals.png

Here, for the exact same time, is a set of time series for every optic except the SRM.  We can see that none of the signals are saturating, and I don't see any big differences for the ITMs or ETMs in the times that the PRMI is locked with high arm powers (center of the x-axis on the plot) and times that the PRMI is not locked, so we don't have high arm powers (edges of the plot - first half second, and last full second).  You can definitely see that the PRM moves much more when the PRMI is locked though, in both pitch and yaw. 

PRMI_2arms_29Oct2013_OpLevs_highArmPowers.png

DCPD signals at the same time:

PRMI_2arms_29Oct2013_DCPDs.png

NB:  These latest 3 plots were created with the getdata script, with arguments "-s 1067163405 -d 7".  It may be a good idea to take some spectra starting at, say 1067163406, 1 second in, and going for ~2 seconds. (It turns out that this is kind of a pain, and I can't convince DTT to give me a sensible spectrum of very short duration....we'll just need to do this live next time around).


Things to think about and investigate:

Why are we losing lock? 

On paper, is the (will the) optical spring a problem once we get high resonance in the arms? 

Spectra of oplevs when we're resonating high arm power.

What is the coupling between 110MHz and 165MHz on the REFL165 PD?  Do we need a stronger bandpass? 

Why are things so shakey when the arm power builds up?

Why do PRCL and MICH have different UGFs when the arms are controlled by ALS vs. ETMs misaligned?

Does QPD for arm transmissions switching work?  Can we then start using TRX and TRY for control?

What is the meaning of the similar features in both transmission signals, and the power recycling cavity?  Power fluctuation in the PRC due to PRM motion? 

  9337   Mon Nov 4 14:11:23 2013 GabrieleSummaryLSCSimulation of REFL_3f signal when the arms come in

Quote:

Yes, the resonance of the 2nd-order sidebands to the IFO screws up the 3f scheme.

2f (~22MHz) and 10f (~110MHz) are at x 5.6 and x 27.9 FSR from the carrier, so that's not the case.

Could we also see how much gain fluctuation of the 3f signals we would experience when the arm comes into the resonance?

 From the simulation there is no visible change in the gain.

  9336   Mon Nov 4 12:59:43 2013 JenneUpdateIOOMC alignment not so good after PSL output shutter installed

Quote:

  The PSL shutter is reinstalled.

 I'm not sure if Steve bumped something, or if it was just a fluke, but the MC didn't come back very nicely after Steve finished re-installing the shutter.

Earlier today, after Steve locked the PMC, MC trans looked good for over an hour (according to the striptool plot on the wall).  Then, the MC was unlocked for about an hour, presumably while Steve was working, he had the light blocked.  When he finished, the MC transmission was around 5,000 while usually it is around 17,000.  The reflection was around 3.4, rather than a best of below 0.5 (unlocked refl is 4.5).

Using Rana's ezcaservo trick to get the suspensions back to where they were at last good lock usually works (I used to do it by hand though).  However, today, it only got the reflection down to about 2.0.  So, I did the rest of the alignment by hand. 

After I did this, the reflection is down to 0.48.  Engaging the WFS makes the MC much more noisy, so I have them disabled currently. 

I have measured the spots, and if I compare them to the measurements that (I think it was Manasa) took last week, they look pretty bad. 

I think that we need to swap out the 2nd zigzag mirror, and then do a careful MC realignment.  It's certainly not worth doing the work, and then re-doing it after we swap out the zigzag mirror.

MCspots_4Nov2013.png

  9335   Mon Nov 4 12:07:55 2013 DmassOmnistructureGeneralReplaced Broken Drill Bits

I broke a small bit while using the 40m drill press to vent some 1/4-20 screws for the cryo experiment.

I replaced it and refilled the small bit row in the bit index I was using; there were ~10 missing sizes

  9334   Mon Nov 4 11:37:12 2013 SteveUpdateIOOPSL output shutter installed

 

 The PSL shutter is reinstalled.The base plate is delrin for isolation and the mount height is adjustable. The last steering mirror mount to be swapped in is ready. It is sitting on the top of the ITMX optical table cover with SS dogs.

There are two reflected spots on the north side of the Uniblitz shutter. They are coming from the vacuum window. They should be trapped also.

Attachment 1: Shutterisback.jpg
Shutterisback.jpg
  9333   Mon Nov 4 09:03:21 2013 SteveUpdatePSLPMC auto locker

Quote:

 I was working on the electronics bench and what sounded like a huge truck rolled by outside. I didn't notice anything until now, but It looks like something became misaligned when the truck passed by (~6:45-6:50 pm). I can hear a lot of noise coming out of the control room speakers and pretty much all of the IOO plots on the wall have sharp discontinuities.

I haven't been moving around much for the past 2 hours so I don't think it was me, but I thought it was worth noting.

 The PMC auto locker is not set to acquire error message made me lock PMC manually.  Arms  locked instantly: TRY 0.9 V and TRX 0.65 V

Attachment 1: 3dTREND.png
3dTREND.png
  9332   Sun Nov 3 00:05:52 2013 CharlesSummaryISSISS Update - Bout' time

Right near the end of summer, I had an ISS board that was nominally working, but had a few problems I couldn't really sort out. Since I've been back, I've spent a lot of time just replacing parts, trying different circuit topologies and generally attempting to make the board function as I hoped it might in all those design stages. Below is a brief list of some of the problems I've been fixing as well as the first good characterization of the board transfer function that I've been able to get.

We'll start with some of the simple problems and proceed to more complicated ones.

  • The 5V reference I was using to obtain an error signal from some arbitrary DC photodiode readout was only producing ~2.5 V. 
    • Turns out I just need a FET type op-amp for the Sallen-Key Filter that I was using to clean up any noise in the reference output, as the leakage current in a AD829 was causing a significant voltage drop. I put in an OPA140 and everything worked marvelously.
  • The way I set up input grounding (i.e. send a ~0 amplitude signal through the board as an input) passed a few Amps through one of my chips causing it to burn out rather fantastically.
    • There isn't a good way to fix this on the current board (besides just getting rid of the functionality altogether) so my solution so far has just been to redesign that particular sub-system/feature and when we implement the second version of the ISS, the input grounding will be done correctly
  • One of the ICs I'm using, specifically the AD8436 RMS-to-DC converter, causes some super strange oscillations in -5V power line. When this chip is soldered onto the board, the -5V supply jumps between -3V and -10V rather sporadically and the DC power-supply used to provide that -5V says that board is drawing ~600 mA on that particular power line.
    • To date, I don't really have any idea what's going with this chip, and I've tried a lot of things to remedy the problem. My first thought was that I had some sort of short somewhere so I took the chip off the board, cleaned up all the excess solder and flux around the chip's footprint and then meticulously soldered a new chip on (when I say meticulously, it took over an hour to solder 20 little feet. I really really didn't want to short anything accidentally as the chip only comes in a package with ridicously small spacing between the leads). Lo and behold, nothing happened. I still saw the same oscillations in power supply and the board was still drawing between >500 mA on that line. Just to be sure, I soldered on a third chip taking the same amount of care and had the same problems.
    • I went over the schematic in Altium that we used to order the board, and unless the manufacturer made a mistake somewhere, there aren't any incorrectly routed signals would cause, say, two active devices to try setting the voltage of a particular node to different values.
    • I got some QSOP-to-DIP package converters so that I could mess around with the AD8436 on a breadboard to make sure it functioned correctly. I set up an identical circuit to the one on the PCB and didn't see any oscillations in the power supply, both for +-5V and +-15V as the chip can handle both supply voltages. I'm not really sure how to interpret this...
    • I'm still actively trying to figure this particular problem out, but I'm shooting in the dark at this point. 
  • Initial attempts to measure the transfer-function of the board were wrought with failure.
    • I figured out, with Nic's help, that the board needs the 'loop closed' with a significant broadband attenuator (to simulate the plant optics discussed in elog 9331) in order to not have constant railing of the high gain op-amp filter stages. Even after I did this, the measured transfer functions were not at all consistent with simulation. I wasn't sure if it was just a part issue, a design issue or a misunderstanding/bad data collection on my part so I just redesigned the whole servo and stuffed the board with entirely new components from around the 40m. Turns out the newly designed servo behaved more properly, as I will show below.

The above list encompasses all the issues I've had in making the ISS board function correctly. No other major problems exist to my knowledge.

I was able to measure both the open- and closed-loop transfer functions of the servo with the SR785. The results are shown below.

full-op-loop.png

The transfer function with the boosts on caps at a particular value set by op-amp railing, i.e. below 100 Hz, the op-amps are already putting out their max voltage. This is the usual physical limitation when measuring the transfer function of an integrator. We can also see that the measured phase follows the simulated phase above ~300 Hz. The 'phase matching' at low frequency is again do to the op-amp railing in the servo output..

The closed-loop gain is shown below,

full-cl-loop.png

The measured closed-loop gain with the boosts on again matches the LISO simulation quite well except at low frequency where we are limited by op-amp railing. We compare the measured closed-loop transfer function to the desired noise suppression stipulated in my previous elog 9331,

req-vs-meas.png

 And we might hopefully conclude that my servo functions as desired. One should note that the op-amp railing seen in these measurements is not indicative of limitations we might face in some application of the ISS for the following reason. These transfer functions were measured with a 100 mV excitation signal (it is necessary to keep this signal amplitude large enough so that the inherent signal-to-noise ratio of the excitation source is large enough for accurate measurement) which leads to somewhat prompt railing of the op-amps. When the ISS operates to actually stabilize a laser, the input error signal will be much smaller (on the order of a few 10's of mV or less) and will decrease significantly assuming correct operation of the ISS. This means we won't see the same type of gain limitations.

 

What now, you ask?

Aside from the problem with the AD8436 chip, the ISS board seems to be functioning correctly. The transfer functions we have measured are correct to within the component tolerances and all of the various subsystems are behaving as they were designed to. Moving toward the goal of having this system work in situ for the CTN experiment, I need to do the following things,

  • Design a housing for the board -> order said housing and the front panel previously designed
  • Make sure the power supply daughter PCB boards are compatible with the ISS board and can provide power correctly
  • Talk to Evan and Tara about integrating the ISS with their experiment and make sure my board can do everything it needs to in that context.

So close, or so I say all the time 

 

  9331   Sat Nov 2 22:49:44 2013 CharlesUpdateISSCTN ISS Noise Suppression Requirement - Updated 10/27

 Previously in elog 8959, I gave a very simple method for determining the noise suppression behavior of the ISS. Recently, I recalculated this requirement in a more correct fashion and again redesigned the ISS to be used in the CTN experiment.

  • Determining the Requirement

Just as before, the data from PSL elog 1270 is necessary to infer a noise suppression requirement. The data presented there by Evan consists of two noise spectra, 1) the unstabilized RIN presently observed in the CTN experiment readout and 2) the theoretical brownian noise produced by thermal processes in the mirror coating+substrate. The statement "TF_mag = (Unstabilized RIN) / (Calculated Brownian Noise Limit)", where TF_mag refers to the required open-loop gain of the ISS, is actually a first order approximation of the 'required' noise suppression. In fact if we wanted the laser noise to be suppressed below the calculated brownian noise level, it is more correct to say 

        Closed-loop ISS gain = (Calculated Brownian Noise Limit) / (Unstabilized RIN)

As this essentially gives a noise suppression spectrum i.e. a closed-loop gain in linear control theory. Below is a very simple block diagram showing how the ISS fits into the CTN experiment. The F(f) block represents my full servo board.

    ISS_path.png

Some of the relevant quantities involved:

            plant-quant_1.png

            plant-quant_2.png

So looking at the block diagram, our full closed-loop transfer function is given by,

cl-loop.png

So then to determine the required F(f), i.e. the required transfer function for my servo, we consider the expression 

               requirement.png

The plant transfer function is simply Plant = (C(f) * a * P * A) ~ 0.014 V/V, where I have ignored the cavity pole around 97 kHz as our open-loop transfer function ends up crossing unity gain around 10 kHz. In the above, I have included what I call a 'safety factor' of 10. Essentially, I want to design my servo such that it suppresses noise well beyond what is actually required so that we can be sure noise contributions to experiment readouts are not significantly influenced by the laser intensity noise.

  • Proposed Servo Design

Using the data Evan reported for the brownian noise and free-running RIN, I came up with an F(f) to the meet the requirement as shown below.

CTN_TF_req-vs-proposed.png

 Where the blue curve includes the safety factor mentioned before. This plot just demonstrates that using my modular ISS design, I can meet the given noise suppression requirements.

To be complete, I'll say a little more about the final design.  As usual, the servo consists of three stages. The first is the usual LP filter that is always 'on' when the ISS loop is closed. The boosts I have chosen to use consist of an integrator with a single zero and a filter that looks somewhat like a de-whitening filter. The simulated open-loop transfer functions are shown below.

switching-filters.png

 

 

 

 

 

 

 

 

  9330   Sat Nov 2 19:36:15 2013 CharlesUpdateGeneralPossible misalignment?

 I was working on the electronics bench and what sounded like a huge truck rolled by outside. I didn't notice anything until now, but It looks like something became misaligned when the truck passed by (~6:45-6:50 pm). I can hear a lot of noise coming out of the control room speakers and pretty much all of the IOO plots on the wall have sharp discontinuities.

I haven't been moving around much for the past 2 hours so I don't think it was me, but I thought it was worth noting.

  9329   Fri Nov 1 19:09:01 2013 rana, evanConfigurationPSLPMC reflected beam nonsense

 While looking at the PMC REFL beam for the AOM diffracted beam, we noticed that although only one beam exists between the PMC and the first steering mirror, there are two afterwards and they both go to the PMC REFL  RFPD!!! This is madness. We only want one beam on our PDH diode.

The reason that we have two beams is that that first steering mirrors is actually a (W1-PW-1025-UV-1064-45P) non-wedged window with an AR coating on only one side. So two beams come out of it. There is a terrible and floppy and illegal anodized aluminum dump close to this beam which *someone* probably intended to use as a "scraper" to get rid of one of the beams.

Black anodized aluminum is a horrible beam dump material at 1064 - its about as grey as Steve's chair. And its so soft that it scatters light back into the PMC and makes more acoustic noise. And it is mounted so poorly (only one screw) that it can easily be bumped and twist and miss the beam. Punchline: only use anodized aluminum dumps for stray light around cameras or for HeNe for OL. Its NOT allowed anywhere where we care about interferometry of NIR beams.

It was also set to dump the dimmer beam. On Monday, we should order ~5 W1 and get them with a wedge of 1-2 deg. Then we use a black glass dump for the dim beam and orient the bright one to hit the REFL camera and the PMC REFL PD.

For the weekend, I have adjusted the crappy grey aluminum flapper to catch the bright beam so that the PMC REFL image no longer shows the interference fringe of two beams. Lets see how the PMC drifts over the next 3 days.

  9328   Fri Nov 1 18:59:41 2013 EvanConfigurationISSAOM cabling

[Rana, Nic, Evan]

We did some work today on getting the AOM back up and running so that we can implement an SR560-based ISS.

We've removed the 18 AWG wire that was previously used to power the driver and have replaced it with a 12 AWG twisted pair (black and white, enclosed in a single gray cladding). This pair runs into the PSL rack's 24 V terminal block with a 2 A fuse. We've also replaced the cable connecting the AOM to the driver; it's now RG405.

Also disconnected the power to the old Kalmus FSS crystal driver box and turned it off. It was powered illegally. Also disconnected the power connection between the Sorensen and the old ISS AA chassis since it was wired directly without any fuse (which is a code violation). It will stay off until someone uses a proper fuse and wiring to hook it back up.

Attachment 1: aom.jpg
aom.jpg
Attachment 2: aom_driver.jpg
aom_driver.jpg
Attachment 3: aom_driver_power.jpg
aom_driver_power.jpg
Attachment 4: 20131101_170120.jpg
20131101_170120.jpg
  9327   Fri Nov 1 17:44:06 2013 KojiSummaryLSCSimulation of REFL_3f signal when the arms come in

Yes, the resonance of the 2nd-order sidebands to the IFO screws up the 3f scheme.

2f (~22MHz) and 10f (~110MHz) are at x 5.6 and x 27.9 FSR from the carrier, so that's not the case.

Could we also see how much gain fluctuation of the 3f signals we would experience when the arm comes into the resonance?

  9326   Fri Nov 1 17:01:46 2013 GabrieleSummaryLSCSimulation of REFL_3f signal when the arms come in

 I simulated how the 3f signal is affected by the resonance condition of the arms.

To keep it simple, I only simulated a double cavity. The attached plot shows the result. In x there is the arm cavity detuning from resonance (in log scale to show what happens close to the 0 value). In the y axis there is the PRC detuning. So every vertical slice of the upper plot gives a PDH signal for a given arm detuning. The bottom plot shows the power build up inside the arm, which is dominated by the carrier.

refl_3f_I_vs_cavity_tuning_noresf2.png

The 3f signal is not perturbed in any significant way by the arm resonance condition. This is good and what we expected.

However, in this simulation I had to ensure that the 1f sidebands are not perfectly anti-resonant inside the arms. They are indeed quite far away from resonance. If the modulation frequency is chosen in order to make the 1f sidebands exactly ant-resonant, the 2f will be resonant. This screws up the signal: REFL_3f is made of two contributions of equal amplitude, one on the PRC sidebands resonance and the other on the PRC carrier resonance. When the arm tuning goes to zero, these two cancels out and there is no more PDH...

refl_3f_I_vs_cavity_tuning.png

However, this is a limit case, since the frequency show match perfectly. If the modulation frequency is few arm line widths away from perfect anti-resonance, we have no problem.

  9325   Fri Nov 1 09:45:32 2013 SteveUpdateIOObeam dumps to be find

Quote:

Quote:

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 I couldn't find any dumps near the WFS. Koji looked. I looked twice. Maybe they are spooky and absorbing all of the light?

The MC alignment was bad and the WFS were making it drift. Koji aligned the beam into the PMC. I then restored the MC suspensions to where they were 8 days ago (back when the transmission and reflection were good). With the WFS OFF, this gave us a MC trans ~ 16000. With WFS ON it goes to 17500 which is about as good as its been over the last 80 days.

I centered the beam on the WFS with the MC unlocked and also centered the beam on the whole WFS path (it was near clipping between WFS 1 & 2). Also for some reason that beamsplitter which steers the beam onto WFS1 is a R=33% (!? why is this not a R=50% ??).

Steve, please swap this out to a BS1-1064-50-1025-45S if we have one sitting around. If not, we want to add this to the CVI purchase list, but not buy until we get a bigger list together.

I also centered this newly aligned beam into the IMC onto the PSL QPDs. We should now use these as a pointing reference for the beam into the IMC.

While doing this I noticed that the beam was almost clipping on the Uniblitz shutter used to block the PSL beam. That shutter is mounted too short and was also not centered horizontally. I removed it for now so that Steve can find a more adjustable mount for it and put it back into play. The beam going into the IMC is BIG, so you have to very careful when centering the shutter. Might be that we cannot leave it at 45 deg and still get a big enough aperture.

Note #3 for Steve: please also replace the mount for last steering mirror into the IMC with a Polanski or a Superman, that black Ultima is no good. Also the dogs must be steel - no aluminum dogs for our sensitive places.

No wonder they could not find the beam dumps. Last night was Haloween. They should of just said: Trick or treat! where are the beam dumps?

Attachment 1: razor_beam_dumps.jpg
razor_beam_dumps.jpg
  9324   Thu Oct 31 21:22:00 2013 rana, kojiSummaryIOOmodulation beat note in MC servo

I hooked up the 4395 to the MC servo board test out (TP2A) and looked at the spectrum using our new SPAG4395.py script. We noticed a huge peak at ~3.8 MHz and correctly guessed that it was due to the beat between the MC modulation frequency 29.5 MHz and 3*f1 (~33 MHz).

So we tuned the Marconi for the main mod. from 11065910 to 11066099 Hz and saw the beat note disappear (to within the 1 Hz tuning precision of our Marconi).

New MC length tuning method! Alert the LA Times!

13031.png13031_200.png13031_200b.png

My conjecture is that this temperature dependent mismatch between the modulation frequency (f1) and the MC length  is what leads sometimes to our nasty saturating PC DRIVE signal. TBD.

  9323   Thu Oct 31 20:05:48 2013 RANAUpdateIOOMode Cleaner Tune-UP

Quote:

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 I couldn't find any dumps near the WFS. Koji looked. I looked twice. Maybe they are spooky and absorbing all of the light?

The MC alignment was bad and the WFS were making it drift. Koji aligned the beam into the PMC. I then restored the MC suspensions to where they were 8 days ago (back when the transmission and reflection were good). With the WFS OFF, this gave us a MC trans ~ 16000. With WFS ON it goes to 17500 which is about as good as its been over the last 80 days.

I centered the beam on the WFS with the MC unlocked and also centered the beam on the whole WFS path (it was near clipping between WFS 1 & 2). Also for some reason that beamsplitter which steers the beam onto WFS1 is a R=33% (!? why is this not a R=50% ??).

Steve, please swap this out to a BS1-1064-50-1025-45S if we have one sitting around. If not, we want to add this to the CVI purchase list, but not buy until we get a bigger list together.

I also centered this newly aligned beam into the IMC onto the PSL QPDs. We should now use these as a pointing reference for the beam into the IMC.

While doing this I noticed that the beam was almost clipping on the Uniblitz shutter used to block the PSL beam. That shutter is mounted too short and was also not centered horizontally. I removed it for now so that Steve can find a more adjustable mount for it and put it back into play. The beam going into the IMC is BIG, so you have to very careful when centering the shutter. Might be that we cannot leave it at 45 deg and still get a big enough aperture.

Note #3 for Steve: please also replace the mount for last steering mirror into the IMC with a Polanski or a Superman, that black Ultima is no good. Also the dogs must be steel - no aluminum dogs for our sensitive places.

Attachment 1: drifty.png
drifty.png
  9322   Thu Oct 31 15:34:28 2013 manasaUpdateIOOMC not happy since yesterday

Quote:

 

8 day minute trend of some of the IMC alignment signals.

That step ~2 days ago in the WFS2 yaw control signal shows that I didn't do such a good job on yaw.

Nic is going to come over some time and give us a new Gouy telescope that let's us have bigger beams on the WFS. At LLO, Hartmut demonstrated recently how bigger beams can reduce offsets somehow...mechanism TBD.

Also, we must angle the WFS and figure out how to dump the reflections at the same time that we rework the table for the telescope.

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 

The MC has not been able to hold lock for over a couple of hours since yesterday. I aligned the MC yesterday (elog 9320) and it lost lock in a couple of hours. I realigned the MC again around noon today only to see it drifting and lose lock again.

I have attached the MC trend for the 2 hours when the MC drifted slowly from its happy to sad state.

 

 

Attachment 1: MC_drift.png
MC_drift.png
ELOG V3.1.3-